OpenStudy (anonymous):

The probability of Keith winning a chess game against Sam is 0.4. What is the probability that Keith will not win 3 games in a row?

jimthompson5910 (jim_thompson5910):

What is the probability of him losing 1 game?

OpenStudy (anonymous):

Are you asking me? cuz I have no idea.... I hate Probability

OpenStudy (anonymous):

But I guess it has to add up to 1 so 1-0.4=0.6 right...??

jimthompson5910 (jim_thompson5910):

good, the probability of him losing any given game is 0.6 so the probability of losing 3 in a row is 0.6*0.6*0.6 = 0.216 shortcut: 0.6^3 = 0.216

OpenStudy (anonymous):

oh that makes sense. I wish my teacher could make sense as well. Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome